An unknown compound was discovered in a laboratory by a chemist. To be able to

find out the molecular formula of this compound, 1.13 g of the unknown compound was

placed into a combustion chamber, and end of the analysis 1.252 g CO2 and 0.3404 g

H2O were obtained. Later, the chemist analyzed the molar mass of the compound as 240

g/mol. What is the molecular formula of the unknown compound?


1
Expert's answer
2021-12-15T01:32:07-0500

Assuming the compound contains C, H and O

And taking molar masses as:

"CO_2=44.011\\H_2O=18.016\\C=12.011\\H=1.008\\O=15.999"


Mass of Carbon in the compound:


"= \frac{1.252}{44.011}×12.011=0.34168"


Mass of Hydrogen in the compound:


"= \frac{0.3404}{18.016}×2×1.008=0.03809g"


Mass of Oxygen in the compound:


"1.13-(0.34168+0.03809)\\=0.75023g"


Moles of "C=\frac{0.34168}{12.011}=0.02845"


Moles of "H=\frac{0.03809}{1.008}=0.03779"



Moles of O"=\frac{0.75023}{15.999}=0.04689"


Simplest ratio:


"\frac{0.02845}{0.02845}:\frac{0.03779}{0.02845}:\frac{0.04689}{0.02845}"


"=1:1.3283:1.64815\\=3(1:1.3283:1.64815)\\\approx3:4:5"


Empirical formula "=C_3H_4O_5"


Mass ="120g"


Molecular formula "\frac{240}{120} =2" times Empirical formula


Molecular formula is:


"C_6H_8O_{10}"